# EraseDiff: Erasing Data Influence in Diffusion Models

### 11 Jan 2024 v1

@@ Abstract
In response to data protection regulations and the “right to be forgotten”, in this work, we introduce an unlearning algorithm for diffusion models. Our algorithm equips a diffusion model with a mechanism to mitigate the concerns related to data memorization. To achieve this, we formulate the unlearning problem as a bi-level optimization problem, wherein the outer objective is to preserve the utility of the diffusion model on the remaining data. The inner objective aims to scrub the information associated with forgetting data by deviating the learnable generative process from the ground-truth denoising procedure. To solve the resulting bi-level problem, we adopt a first-order method, having superior practical performance while being vigilant about the diffusion process and solving a bi-level problem therein. Empirically, we demonstrate that our algorithm can preserve the model utility, effectiveness, and efficiency while removing across two widely-used diffusion models and in both conditional and unconditional image generation scenarios. In our experiments, we demonstrate the unlearning of classes, attributes, and even a race from face and object datasets such as UTKFace, CelebA, CelebA-HQ, and CIFAR10.
@@ 1 Introduction
Diffusion Models (Ho et al., 2020; Song et al., 2020; Rombach et al., 2022) are now the method of choice in deep generative models, owing to their high-quality output, stability, and ease of training procedure. This has facilitated their successful integration into commercial applications such as midjourney [^1] https://docs.midjourney.com/. Unfortunately, the ease of use associated with diffusion models brings forth significant privacy risks. Studies have shown that these models can memorize and regenerate individual images from their training datasets (Somepalli et al., 2023a; b; Carlini et al., 2023). Beyond privacy, diffusion models are susceptible to misuse, capable of generating inappropriate digital content (Rando et al., 2022; Salman et al., 2023; Schramowski et al., 2023). They are also vulnerable to poison attacks (Chen et al., 2023b), allowing the generation of target images with specific triggers. These factors collectively pose substantial security threats. Moreover, the ability of diffusion models to emulate distinct artistic styles (Shan et al., 2023; Gandikota et al., 2023a) raises questions about data ownership and compliance with intellectual property and copyright laws.
In this context, individuals whose images are used for training might request the removal of their private data. In particular, data protection regulations like the European Union General Data Protection Regulation (GDPR) (Voigt & Von dem Bussche, 2017) and the California Consumer Privacy Act (CCPA) (Goldman, 2020) grant users the right to be forgotten, obligating companies to expunge data pertaining to a user upon receiving a request for deletion. These legal provisions grant data owners the right to remove their data from trained models and eliminate its influence on said models (Bourtoule et al., 2021; Guo et al., 2020; Golatkar et al., 2020; Mehta et al., 2022; Sekhari et al., 2021; Ye et al., 2022; Tarun et al., 2023b; a; Chen et al., 2023a).
@@
In this work, we propose a method to scrub the data information from the diffusion models without requiring training the whole system from scratch. Specifically, the proposed method EraseDiff formulates diffusion unlearning as a bi-level optimization problem, where the outer objective is to finetune the models with the remaining data for preserving the model utility and the inner objective aims to erase the influence of the forgetting data on the models by deviating the learnable reverse process from the ground-truth denoising procedure. Then, a first-order solution is adopted to solve the resulting problem. We benchmark EraseDiff on various scenarios, encompassing unlearning of classes on CIFAR10 (Krizhevsky et al., 2009) and races on UTKFace (Zhang et al., 2017) with conditional diffusion models, attributes on CelebA (Liu et al., 2015) as well as CelebA-HQ (Lee et al., 2020) with unconditional diffusion models. The results demonstrate that EraseDiff surpasses the baseline methods for diffusion unlearning across a range of evaluation metrics.
@@ 2 Background
In the following section, we outline the components of the models we evaluate, including Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020), denoising diffusion implicit models (DDIM) (Song et al., 2020), and classifier-free guidance diffusion models (Ho & Salimans, 2022). Throughout the paper, we denote scalars, and vectors/matrices by lowercase and bold symbols, respectively (e.g., $a$, ${\bm{a}}$, and ${\bm{A}}$).
@@
Classifier-free guidance is a conditioning method to guide diffusion-based generative models without an external pre-trained classifier. Model prediction would be $\epsilon_{{\bm{\theta}}}({\mathbf{x}}_{t},t,c)$, where $c$ is the input’s corresponding label. The unconditional and conditional models are jointly trained by randomly setting $c$ to the unconditional class $\emptyset$ identifier with the probability $p_{uncond}$. Then, the sampling procedure would use the linear combination of the conditional and unconditional score estimates as ${\bm{\epsilon}}_{t}=(1+w)\cdot\epsilon_{{\bm{\theta}}}({\mathbf{x}}_{t},t,c)-w\cdot\epsilon_{{\bm{\theta}}}({\mathbf{x}}_{t},t)$, and $w$ is the guidance scale that controls the strength of the classifier guidance.
@@ 3.1 Training objective
Our goal is to scrub the information about the forgetting data ${\mathcal{D}}_{f}$ carried by the diffusion models while maintaining the model utility over the remaining data ${\mathcal{D}}_{r}$. To achieve this, we adopt different training objectives for ${\mathcal{D}}_{r}$ and ${\mathcal{D}}_{f}$ as follows.
@@
For the remaining data ${\mathcal{D}}_{r}$, we finetune the diffusion models with the original objective by minimizing the variational bound on negative log-likelihood:
$\min_{{\bm{\theta}}}\mathbb{E}_{{\mathbf{x}}_{0}\sim{\mathcal{D}}_{r}}[-\log p_{\bm{\theta}}({\mathbf{x}}_{0})]\propto\min_{{\bm{\theta}}}\mathbb{E}_{{\mathbf{x}}_{0}\sim{\mathcal{D}}_{r}}\bigg{[}\sum_{t=2}^{T}\underbrace{\text{KL}(q({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t},{\mathbf{x}}_{0})\|p_{{\bm{\theta}}}({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t}))}_{{\mathcal{L}}_{t-1}({\bm{\theta}},{\mathcal{D}}_{r})}\bigg{]},$ (1)
where ${\mathbf{x}}_{t}=\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}$ with ${\bm{\epsilon}}\in{\mathcal{N}}({\bm{0}},\mathbf{I}_{d})$. Given a large $T$ and if $\alpha_{T}$ is sufficiently close to 0, $x_{T}$ would converge to a standard Gaussian distribution, so $p_{\theta}({\mathbf{x}}_{T}):={\mathcal{N}}({\bm{0}},\mathbf{I}_{d})$. Eq. 1 aims to minimize the KL divergence between the ground-truth backward distribution $q({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t},{\mathbf{x}}_{0})={\mathcal{N}}({\mathbf{x}}_{t-1};\tilde{{\bm{\mu}}}_{t},\tilde{\beta}_{t}^{2}\mathbf{I}_{d})$ and the learnable backward distribution $p_{{\bm{\theta}}}({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})$. With $p_{{\bm{\theta}}}({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})={\mathcal{N}}({\mathbf{x}}_{t-1};{\bm{\mu}}_{\bm{\theta}}({\mathbf{x}}_{t},t),\sigma_{t}^{2}\mathbf{I}_{d})$, then we define
${\mathcal{F}}({\bm{\theta}}):={\mathcal{L}}_{t-1}({\bm{\theta}},{\mathcal{D}}_{r})=\mathbb{E}_{{\mathbf{x}}_{0}\sim{\mathcal{D}}_{r},{\bm{\epsilon}}\sim{\mathcal{N}}({\bm{0}},\mathbf{I}_{d})}\Big{[}a\cdot\big{\|}{\bm{\epsilon}}-\epsilon_{{\bm{\theta}}}(\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}},t)\big{\|}^{2}\Big{]},$ (2)
where the coefficient $a=\frac{\beta_{t}^{2}}{2\sigma_{t}^{2}\alpha_{t}(1-\bar{\alpha}_{t})}$ for DDPM and $a=\frac{(\sqrt{\alpha_{t}(1-\bar{\alpha}_{t-1}-\sigma_{t}^{2})}-\sqrt{1-\bar{\alpha}_{t}})^{2}}{2\sigma_{t}^{2}\alpha_{t}}$ for DDIM. Eq. 2 constraints the model $\epsilon_{\bm{\theta}}$ to predict ${\bm{\epsilon}}$ from ${\mathbf{x}}_{t}$, with the goal of aligning the learnable backward distribution $p_{{\bm{\theta}}}({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})$ closely with the ground-truth backward distribution $q({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t},{\mathbf{x}}_{0})$.
@@
$\max_{{\hat{{\bm{\theta}}}}}\mathbb{E}_{{\mathbf{x}}_{0}\sim{\mathcal{D}}_{f}}[-\log p_{{\hat{{\bm{\theta}}}}}({\mathbf{x}}_{0})]\propto\max_{{\hat{{\bm{\theta}}}}}\mathbb{E}_{{\mathbf{x}}_{0}\sim{\mathcal{D}}_{f}}\bigg{[}\sum_{t=2}^{T}\underbrace{\text{KL}(q({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t},{\mathbf{x}}_{0})\|p_{{{\hat{{\bm{\theta}}}}}}({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t}))}_{{\mathcal{L}}_{t-1}(\hat{{\bm{\theta}}},{\mathcal{D}}_{f})}\bigg{]}.$ (3)
@@
$\hat{{\bm{\epsilon}}}_{t}$ could be any distribution different from ${\bm{\epsilon}}_{t}$, we choose uniform distribution $\hat{{\bm{\epsilon}}}_{t}\sim{\mathcal{U}}({\bm{0}},{\bm{I}})$ for experiments due to no extra hyper-parameters being needed. Appendix A.2 also show results for $\hat{{\bm{\epsilon}}}_{t}\sim{\mathcal{N}}({\bm{\mu}},\mathbf{I}_{d})$ where ${\bm{\mu}}\neq{\bm{0}}$.
Then, we define the following objective function
@@
$f(\hat{{\bm{\theta}}},{\mathcal{D}}_{f}):=\mathbb{E}_{{\mathbf{x}}_{0}\sim{\mathcal{D}}_{f},{\bm{\epsilon}}\sim{\mathcal{N}}({\bm{0}},{\bm{I}}),\hat{{\bm{\epsilon}}}\sim{\mathcal{U}}({\bm{0}},{\bm{1}})}\Big{[}a\cdot\big{\|}\hat{{\bm{\epsilon}}}-\epsilon_{{{\hat{{\bm{\theta}}}}}}(\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}},t)\big{\|}^{2}\Big{]}\;.$ (4)
@@
With this, the scrubbed model $\epsilon_{\hat{{\bm{\theta}}}}$ would tend to predict $\hat{{\bm{\epsilon}}}$ given ${\mathbf{x}}_{t}$. As such, for the forgetting data, the approximator $\epsilon_{\hat{{\bm{\theta}}}}$ cannot learn the correct denoising distribution and thus cannot help to generate corresponding images when sampling.
Final objective.
We strive to learn a model $\epsilon_{{\bm{\theta}}}$, such that when we apply the unlearning algorithm $\text{Alg}({\bm{\theta}},{\mathcal{D}}_{f})$ to erase the influence of the forgetting data ${\mathcal{D}}_{f}$ in the model, the resulting model with parameters ${\bm{\theta}}^{*}$ should still be adept at generating samples having the same distribution as ${\mathcal{D}}_{r}$. To achieve this, given a well-trained diffusion model with parameters ${\bm{\theta}}_{0}$ over the data ${\mathcal{D}}$, for $t={2,\cdots,T}$, refer to  Rajeswaran et al. (2019), we update the model with the following objective
${\bm{\theta}}^{*}:=\operatorname*{arg\,min}_{{\bm{\theta}}}{\mathcal{F}}({\bm{\theta}}),\hskip 10.00002pt\text{where}\hskip 5.0pt{\mathcal{F}}({\bm{\theta}})={\mathcal{L}}(\text{Alg}({\bm{\theta}},{\mathcal{D}}_{f}),{\mathcal{D}}_{r}),$ (5)
where ${\mathbb{R}}^{d}\ni\bm{\phi}^{*}({\bm{\theta}}):=\text{Alg}({\bm{\theta}},{\mathcal{D}}_{f})=\operatorname*{arg\,min}_{\bm{\phi}}f(\bm{\phi},{\mathcal{D}}_{f})$, $\bm{\phi}$ is initialized with ${\bm{\theta}}$ and ${\bm{\theta}}$ starts from ${\bm{\theta}}_{0}$. Starting from the parameters ${\bm{\theta}}_{0}$, for incoming forgetting data ${\mathcal{D}}_{f}$, Eq. 5 updates the approximator $\epsilon_{{\bm{\theta}}_{0}}$ to scrub the information about ${\mathcal{D}}_{f}$, and the updated weights ${\bm{\theta}}^{*}$ can also preserve the model utility over the remaining data ${\mathcal{D}}_{r}$. To encompass the unconditional and conditional image generation scenarios, Eq. 5 can be further rewritten as
$\min_{{\bm{\theta}}}\mathbb{E}_{{\mathbf{x}}_{0}\sim{\mathcal{D}}_{r},c\sim{\mathcal{C}}_{r}}\Big{[}a\cdot\big{\|}{\bm{\epsilon}}-\epsilon_{{\bm{\theta}}}(\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}},t,c)\big{\|}^{2}\Big{]},$
s.t. ${\bm{\theta}}\in\operatorname*{arg\,min}_{\hat{{\bm{\theta}}}}\mathbb{E}_{{\mathbf{x}}_{0}\sim{\mathcal{D}}_{f},\hat{{\bm{\epsilon}}}\sim{\mathcal{U}}({\bm{0}},{\bm{1}}),c\sim{\mathcal{C}}_{f}}\Big{[}a\cdot\big{\|}\hat{{\bm{\epsilon}}}-\epsilon_{{{\hat{{\bm{\theta}}}}}}(\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}},t,c)\big{\|}^{2}\Big{]},$ (6)
where $c$ would be $\emptyset$ for unconditional diffusion models and labels for conditional diffusion models. Eqs. 5 and 3.1 aim to keep the model utility over the remaining data ${\mathcal{D}}_{r}$, but misguide it ${\mathcal{D}}_{f}$ by deviating the learnable reverse process from the ground-truth denoising procedure.
3.2 Solution.

### 05 Feb 2024 v2

@@ Abstract
In this work, we introduce an unlearning algorithm for diffusion models. Our algorithm equips a diffusion model with a mechanism to mitigate the concerns related to data memorization. To achieve this, we formulate the unlearning problem as a constraint optimization problem, aiming to preserve the utility of the diffusion model on the remaining data and scrub the information associated with forgetting data by deviating the learnable generative process from the ground-truth denoising procedure. To solve the resulting problem, we adopt a first-order method, having superior practical performance while being vigilant about the diffusion process. Empirically, we demonstrate that our algorithm can preserve the model utility, effectiveness, and efficiency while removing across the widely-used diffusion models and in both conditional and unconditional image generation scenarios.
Figure 1: Generated samples by our method, EraseDiff, to erase the targeted class/concept. EraseDiff can forget classes and avoid NSFW content.
@@ 1 Introduction
Diffusion Models Ho et al. (2020); Song et al. (2020); Rombach et al. (2022) are now the method of choice in deep generative models, owing to their high-quality output, stability, and ease of training procedure. This has facilitated their successful integration into commercial applications such as midjourney [^1] https://docs.midjourney.com/. Unfortunately, the ease of use associated with diffusion models brings forth significant privacy risks. Studies have shown that these models can memorize and regenerate individual images from their training datasets Somepalli et al. (2023a; b); Carlini et al. (2023). Beyond privacy, diffusion models are susceptible to misuse and can generate inappropriate digital content Rando et al. (2022); Salman et al. (2023); Schramowski et al. (2023). They are also vulnerable to poison attacks Chen et al. (2023b), allowing the generation of target images with specific triggers. These factors collectively pose substantial security threats. Moreover, the ability of diffusion models to emulate distinct artistic styles Shan et al. (2023); Gandikota et al. (2023a) raises questions about data ownership and compliance with intellectual property and copyright laws.
In this context, individuals whose images are used for training might request the removal of their private data. In particular, data protection regulations like the European Union General Data Protection Regulation (GDPR) Voigt & Von dem Bussche (2017) and the California Consumer Privacy Act (CCPA) Goldman (2020) grant users the right to be forgotten, obligating companies to expunge data pertaining to a user upon receiving a request for deletion. These legal provisions grant data owners the right to remove their data from trained models and eliminate its influence on said models Bourtoule et al. (2021); Guo et al. (2020); Golatkar et al. (2020); Mehta et al. (2022); Sekhari et al. (2021); Ye et al. (2022); Tarun et al. (2023b; a); Chen et al. (2023a).
@@
Recently, a handful of studies Gandikota et al. (2023a; b); Heng & Soh (2023a; b); Fan et al. (2023); Zhang et al. (2023) target unlearning in diffusion models. Most of them target text-to-image diffusion models Gandikota et al. (2023a; b); Heng & Soh (2023a); Zhang et al. (2023).  Heng & Soh (2023b) introduce a versatile framework applicable to a wide range of generative models. However, the method proposed by  Heng & Soh (2023b) requires the computation of the Fisher Information Matrix (FIM) for different datasets and models, which may lead to significant computational demands.
In this work, we propose EraseDiff to scrub the data information from the diffusion models without requiring training the whole system from scratch. Specifically, EraseDiff formulates diffusion unlearning as a constraint optimization problem, where the objective is to finetune the models with the remaining data for preserving the model utility and to erase the influence of the forgetting data on the models by deviating the learnable reverse process from the ground-truth denoising procedure. Then, a first-order solution is adopted to solve the problem. We benchmark EraseDiff on various scenarios, encompassing unlearning of classes on CIFAR10 Krizhevsky et al. (2009) with conditional diffusion models, attributes on CelebA-HQ Lee et al. (2020) with unconditional diffusion models, classes on Imagenette Howard & Gugger (2020) and concepts on the I2P dataset Schramowski et al. (2023) with stable diffusion. Our findings reveal that EraseDiff achieves more than twice the speed compared to Heng and Soh’s method Heng & Soh (2023b) in terms of a single training step, note that this comparison does not take into account the additional time required for computing the FIM in Heng and Soh’s method Heng & Soh (2023b). The results demonstrate that EraseDiff is capable of effectively erasing data influence in diffusion models, including a wide spectrum of categories, ranging from specific classes and attributes to nudity.
@@ 2 Background
In this section, we outline the components of the models we evaluate, including Denoising Diffusion Probabilistic Models (DDPM) Ho et al. (2020), denoising diffusion implicit models (DDIM) Song et al. (2020), classifier-free guidance diffusion models Ho & Salimans (2022), and Latent Diffusion Models (LDM) Rombach et al. (2022). Throughout the paper, we denote scalars, and vectors/matrices by lowercase and bold symbols, respectively (e.g., $a$, ${\bm{a}}$, and ${\bm{A}}$).
@@
Classifier-free guidance is a conditioning method to guide diffusion-based generative models without an external pre-trained classifier. Model prediction would be $\epsilon_{{\bm{\theta}}}({\mathbf{x}}_{t},t,c)$, where $c$ is the input’s corresponding label. The unconditional and conditional models are jointly trained by randomly setting $c$ to the unconditional class $\emptyset$ identifier with the probability $p_{uncond}$. Then, the sampling procedure uses the linear combination of the conditional and unconditional score estimates as ${\bm{\epsilon}}_{t}=(1+w)\cdot\epsilon_{{\bm{\theta}}}({\mathbf{x}}_{t},t,c)-w\cdot\epsilon_{{\bm{\theta}}}({\mathbf{x}}_{t},t)$, $w$ is the guidance scale that controls the strength of the classifier guidance.
Latent diffusion model.
Latent diffusion models (LDM) apply the diffusion models in the latent space ${\mathbf{z}}$ of a pre-trained variational autoencoder. The noise would be added to ${\mathbf{z}}=\varepsilon({\mathbf{x}})$, instead of the data ${\mathbf{x}}$, and the denoised output would be transformed to image space with the decoder. Besides, cross-attention layers are introduced into the model for general conditioning inputs.
@@ 3.1 Training objective
Our goal is to scrub the information about ${\mathcal{D}}_{f}$ carried by the diffusion models while maintaining the model utility over the remaining data ${\mathcal{D}}_{r}$. To achieve this, we adopt different training objectives for ${\mathcal{D}}_{r}$ and ${\mathcal{D}}_{f}$ as follows.
@@
For the remaining data ${\mathcal{D}}_{r}$, we fine-tune the diffusion models with the original objective by minimizing the variational bound on negative log-likelihood:
$\text{min}_{{\bm{\theta}}}\mathbb{E}_{{\mathbf{x}}_{0}\sim{\mathcal{D}}_{r}}[-\log p_{\bm{\theta}}({\mathbf{x}}_{0})]\propto\text{min}_{{\bm{\theta}}}\mathbb{E}_{{\mathbf{x}}_{0}\sim{\mathcal{D}}_{r}}\bigg{[}\sum_{t=2}^{T}\underbrace{\text{KL}(q({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t},{\mathbf{x}}_{0})\|p_{{\bm{\theta}}}({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t}))}_{{\mathcal{L}}_{t-1}({\bm{\theta}},{\mathcal{D}}_{r})}\bigg{]},$ (1)
where ${\mathbf{x}}_{t}=\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}$ with ${\bm{\epsilon}}\in{\mathcal{N}}({\bm{0}},\mathbf{I}_{d})$. Given a large $T$ and if $\alpha_{T}$ is sufficiently close to 0, $x_{T}$ would converge to a standard Gaussian distribution, so $p_{\theta}({\mathbf{x}}_{T}):={\mathcal{N}}({\bm{0}},\mathbf{I}_{d})$. Eq. 1 aims to minimize the KL divergence between the ground-truth backward distribution $q({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t},{\mathbf{x}}_{0})={\mathcal{N}}({\mathbf{x}}_{t-1};\tilde{{\bm{\mu}}}_{t},\tilde{\beta}_{t}^{2}\mathbf{I}_{d})$ and the learnable backward distribution $p_{{\bm{\theta}}}({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})$. With $p_{{\bm{\theta}}}({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})={\mathcal{N}}({\mathbf{x}}_{t-1};{\bm{\mu}}_{\bm{\theta}}({\mathbf{x}}_{t},t),\sigma_{t}^{2}\mathbf{I}_{d})$, then we have
${\mathcal{L}}({\bm{\theta}},{\mathcal{D}}_{r}):=\mathbb{E}_{{\mathbf{x}}_{0}\sim{\mathcal{D}}_{r},{\bm{\epsilon}}\sim{\mathcal{N}}({\bm{0}},\mathbf{I}_{d}),t}\Big{[}a_{t}\cdot\big{\|}{\bm{\epsilon}}-\epsilon_{{\bm{\theta}}}(\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}},t)\big{\|}^{2}\Big{]},$ (2)
where the coefficient $a_{t}=\frac{\beta_{t}^{2}}{2\sigma_{t}^{2}\alpha_{t}(1-\bar{\alpha}_{t})}$ for DDPM and $a_{t}=\frac{(\sqrt{\alpha_{t}(1-\bar{\alpha}_{t-1}-\sigma_{t}^{2})}-\sqrt{1-\bar{\alpha}_{t}})^{2}}{2\sigma_{t}^{2}\alpha_{t}}$ for DDIM. Eq. 2 constraints the model $\epsilon_{\bm{\theta}}$ to predict ${\bm{\epsilon}}$ from ${\mathbf{x}}_{t}$, with the goal of aligning the learnable backward distribution $p_{{\bm{\theta}}}({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})$ closely with the ground-truth backward distribution $q({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t},{\mathbf{x}}_{0})$.
@@
$\max_{{\hat{{\bm{\theta}}}}}\mathbb{E}_{{\mathbf{x}}_{0}\sim{\mathcal{D}}_{f}}[-\log p_{{\hat{{\bm{\theta}}}}}({\mathbf{x}}_{0})]\propto\max_{{\hat{{\bm{\theta}}}}}\mathbb{E}_{{\mathbf{x}}_{0}\sim{\mathcal{D}}_{f}}\bigg{[}\sum_{t=2}^{T}\underbrace{\text{KL}(q({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t},{\mathbf{x}}_{0})\|p_{{{\hat{{\bm{\theta}}}}}}({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t}))}_{{\mathcal{L}}_{t-1}(\hat{{\bm{\theta}}},{\mathcal{D}}_{f})}\bigg{]}.$
@@
$\hat{{\bm{\epsilon}}}_{t}$ could be any distribution different from ${\bm{\epsilon}}_{t}$, we choose the uniform distribution $\hat{{\bm{\epsilon}}}_{t}\sim{\mathcal{U}}({\bm{0}},{\bm{I}})$ for experiments on DDPM and DDIM due to no extra hyper-parameters being needed. Appendix also shows results for $\hat{{\bm{\epsilon}}}_{t}\sim{\mathcal{N}}({\bm{\mu}},\mathbf{I}_{d})$ where ${\bm{\mu}}\neq{\bm{0}}$.
Then, we define the following objective function:
@@
$f(\hat{{\bm{\theta}}},{\mathcal{D}}_{f})$ $:=\mathbb{E}_{{\mathbf{x}}_{0}\sim{\mathcal{D}}_{f},{\bm{\epsilon}}\sim{\mathcal{N}}({\bm{0}},{\bm{I}}),\hat{{\bm{\epsilon}}}\sim{\mathcal{U}}({\bm{0}},\mathbf{I}),t}\Big{[}a_{t}\cdot\big{\|}\hat{{\bm{\epsilon}}}-\epsilon_{{{\hat{{\bm{\theta}}}}}}(\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}},t)\big{\|}^{2}\Big{]}\;.$ (3)
@@
With this, the scrubbed model $\epsilon_{\hat{{\bm{\theta}}}}$ would tend to predict $\hat{{\bm{\epsilon}}}$ given ${\mathbf{x}}_{t}$. As such, for ${\mathcal{D}}_{f}$, the approximator $\epsilon_{\hat{{\bm{\theta}}}}$ cannot learn the correct denoising distribution and thus cannot help to generate corresponding images when sampling.
3.2 Final objective and solution
Inspired by Liu et al. Liu et al. (2022), we consider the following optimization problem:
$\operatorname{min}_{{\bm{\theta}}}\;{\mathcal{L}}\left({\bm{\theta}},{\mathcal{D}}_{r}\right)\quad\quad\text{s.t.}\quad f\left({\bm{\theta}},{\mathcal{D}}_{f}\right)-\operatorname{min}_{\bm{\phi}\mid{\bm{\theta}}}f\left(\bm{\phi},{\mathcal{D}}_{f}\right)\leq 0.$ (4)
It is worth noting that the constraints $f\left({\bm{\theta}},{\mathcal{D}}_{f}\right)-\operatorname{min}_{\bm{\phi}\mid{\bm{\theta}}}f\left(\bm{\phi},{\mathcal{D}}_{f}\right)\leq 0$ implies ${\bm{\theta}}\in\operatorname{argmin}_{{\bm{\theta}}^{\prime}}f\left({\bm{\theta}}^{\prime},{\mathcal{D}}_{f}\right)$. This further means that among ${\bm{\theta}}$ that minimizes $f\left({\bm{\theta}},{\mathcal{D}}_{f}\right)$, we select the one that minimizes ${\mathcal{L}}\left({\bm{\theta}},{\mathcal{D}}_{r}\right)$.
Denote $g\left({\bm{\theta}}\right)=f\left({\bm{\theta}},{\mathcal{D}}_{f}\right)-\operatorname{min}_{\bm{\phi}\mid{\bm{\theta}}}f\left(\bm{\phi},{\mathcal{D}}_{f}\right)$. Given the current solution ${\bm{\theta}}_{k}$, we aim to update
${\bm{\theta}}_{k+1}={\bm{\theta}}_{k}-\eta\bm{\delta}_{k},$
so that ${\mathcal{L}}\left({\bm{\theta}}_{k+1},{\mathcal{D}}_{r}\right)$ decreases and $g\left({\bm{\theta}}_{k+1}\right)$ decreases if $g\left({\bm{\theta}}_{k}\right)>0$. To this end, we find $\bm{\delta}_{k}$ by
$\bm{\delta}_{k}\in\frac{1}{2}\operatorname{argmin}_{\bm{\delta}}\big{\|}\nabla_{{\bm{\theta}}}{\mathcal{L}}\left({\bm{\theta}}_{k},{\mathcal{D}}_{r}\right)-\bm{\delta}\big{\|}_{2}^{2}\quad\quad\text{s.t.}\quad\;\nabla_{{\bm{\theta}}}g\left({\bm{\theta}}_{k}\right)^{\top}\bm{\delta}\geq a_{k}>0.$ (5)
Because $g\left({\bm{\theta}}_{k+1}\right)-g\left({\bm{\theta}}_{k}\right)\approx-\eta\nabla_{{\bm{\theta}}}g\left({\bm{\theta}}_{k}\right)^{\top}\bm{\delta}\leq-\eta a_{k}<0$, we can ensure that $g\left({\bm{\theta}}_{k+1}\right)<g\left({\bm{\theta}}_{k}\right)$ for small $\eta>0$. To find the solution to the optimization problem in Eq. (5), we develop the following theorem:
Theorem 1.
The optimal solution of the optimization problem in Eq. (5) is $\bm{\delta}^{*}=\nabla_{{\bm{\theta}}}{\mathcal{L}}\left({\bm{\theta}}_{k},{\mathcal{D}}_{r}\right)+\lambda_{k}\nabla_{{\bm{\theta}}}g\left({\bm{\theta}}_{k}\right)$ where $\lambda_{k}=\max\left\{0,\frac{a_{k}-\nabla_{{\bm{\theta}}}g\left({\bm{\theta}}_{k}\right)^{\top}\nabla_{{\bm{\theta}}}{\mathcal{L}}\left(\bm{\theta_{k}},{\mathcal{D}}_{r}\right)}{\|\nabla_{{\bm{\theta}}}g\left({\bm{\theta}}_{k}\right)\|_{2}^{2}}\right\}.$
Proof.
We construct the Lagrange function with $\lambda\geq 0$:
$h\left(\bm{\delta},\lambda\right)=\frac{1}{2}\big{\|}\nabla_{{\bm{\theta}}}{\mathcal{L}}\left(\bm{\theta_{k}},{\mathcal{D}}_{r}\right)-\bm{\delta}\big{\|}_{2}^{2}+\lambda\left(a_{k}-\nabla_{{\bm{\theta}}}g\left({\bm{\theta}}_{k}\right)^{\top}\bm{\delta}\right).$
Using Karush-Kuhn-Tucker (KKT) theorem, at the optimal solution, we have $\lambda\geq 0$ and
$\bm{\delta}-\nabla_{{\bm{\theta}}}{\mathcal{L}}\left(\bm{\theta_{k}},{\mathcal{D}}_{r}\right)-\lambda\nabla_{{\bm{\theta}}}g\left({\bm{\theta}}_{k}\right)$ $=\bm{0},$
$\nabla_{{\bm{\theta}}}g\left({\bm{\theta}}_{k}\right)^{\top}\bm{\delta}$ $\geq a_{k},$
$\lambda\left(a_{k}-\nabla_{{\bm{\theta}}}g\left({\bm{\theta}}_{k}\right)^{T}\bm{\delta}\right)$ $=0\;.$
From the above constraints, we obtain
$\bm{\delta}$ $=\nabla_{{\bm{\theta}}}{\mathcal{L}}\left(\bm{\theta_{k}},{\mathcal{D}}_{r}\right)+\lambda\nabla_{{\bm{\theta}}}g\left({\bm{\theta}}_{k}\right),\lambda=\max\left\{0,\frac{a_{k}-\nabla_{{\bm{\theta}}}g\left({\bm{\theta}}_{k}\right)^{T}\nabla_{{\bm{\theta}}}{\mathcal{L}}\left(\bm{\theta_{k}},{\mathcal{D}}_{r}\right)}{\|\nabla_{{\bm{\theta}}}g\left({\bm{\theta}}_{k}\right)\|_{2}^{2}}\right\}.$
Moreover, in practice we can choose $a_{k}=\eta\|\nabla_{{\bm{\theta}}}g\left({\bm{\theta}}_{k}\right)\|_{2}^{2}$. The remaining question is how to compute $\nabla_{{\bm{\theta}}}g\left({\bm{\theta}}_{k}\right)$. For this computation, we start from $\bm{\phi}^{0}={\bm{\theta}}_{k}$ and use gradient descend in $K$ steps with the learning rate $\xi>0$ to reach $\bm{\phi}^{K}$. Finally, we compute
$\nabla_{{\bm{\theta}}}g\left({\bm{\theta}}_{k}\right)=\nabla_{{\bm{\theta}}}f\left({\bm{\theta}}_{k},{\mathcal{D}}_{f}\right)-\nabla_{{\bm{\theta}}}f\left(\bm{\phi}^{k},{\mathcal{D}}_{f}\right).$